Calculus I, Fall 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 09

1. Section 4.7: problem 9. Then in problems 18, 21, 29, 30, give formula for Newton’s method, and need not find the numerical value of the solution.

In problem 29, also estimate how many iterations it takes to get |x_{n} − x_{∗}| ≤ 10^{−3}.

Hint for problem 18: you need to find an equation so that π is a solution. Trigonometric functions will help.

2. Section 4.8: pick at least 10 among the odd ones in problems 25-69. Problems 79, 80, 84, 127, 128.